Given the function h described by h(x) = 5x, find each of the following:
(a) h (-14)=
(b) h (20) =
(c) h (30) =

Answer by likaaka(51)

Just substitute in the value of x and multiply.
Given the function h described by h(x) = 5x, find each of the following:
(a) h (-14)= 5(-14) = -70
(b) h (20) = 5(20) = 100
(c) h (30) = 5(30) = 150
The way I remember this is that functions are equations of lines and h(x) is y so when x = -14, then y = -70 and (-14, -70) is a point on the line of y = 5x.
(20, 100) and (30, 150) are also points on the line y = 5x